Checkers uses the periodic inventory system. For the current month, the beginning inventory consisted of 7,200 units that cost $12 each. During the month, the company made two purchases: 3,000 units at $13 each as the first purchase and 12,000 units at $13.50 each as the second purchase. Checkers also sold 12,900 units during the month. Using the average cost method, what is the amount of cost of goods sold for the month?

a.   $167,055.

b.   $173,700.

c.   $161,850.

d.   None of above
Answer #1

Total cost of goods available for sale

= (7200*12)+(3000*13)+(12000*13.50) = 287,400

Totsl units = 7200+3000+12000 = 22,200

Average cost per unit = 287,400/22,200 = 12.95 per unit

Cost of Goods Sold = 12900*12.95

= 167,055